OpenStudy (emilyjones284):

Please help for medal and fan and just to be nice (: Suppose you roll two number cubes and find the probability distribution for the sum of the numbers. Which two sums have the same probability distribution and would be represented with equal bars on a bar graph? A. 4 and 11 B. 6 and 9 C. 5 and 10 D. 2 and 12

OpenStudy (dumbcow):

well "7" is most likely, then its a symmetric distribution... 6 same as 8 5 same as 9 and so on...

OpenStudy (emilyjones284):

wait i dont understand...what would the answer be then?

OpenStudy (anonymous):

fill in the table, and you will see the number of ways to get each roll

OpenStudy (emilyjones284):

do i fill it in by multiplying the two dice or adding them?

OpenStudy (anonymous):

add

OpenStudy (anonymous):

as you add the tops of the dice to get each total

OpenStudy (emilyjones284):

and then what do i do after that?

OpenStudy (emilyjones284):

how do i know what the answer is?

OpenStudy (anonymous):

count the number of ways to get each total for example there are 3 ways to get a total of 4, so the probability you get a 4 is \(\frac{3}{36}\)

OpenStudy (anonymous):

you will notice that there is only one way to get a total of 2 (namely (1,1)) and also only one way to get a total of 12 (6,6)

OpenStudy (emilyjones284):

im still confused haha

OpenStudy (emilyjones284):

counts the totals of how to get each number?

OpenStudy (anonymous):

ok lets go slow

OpenStudy (anonymous):

when you fill out the table you will see the number of ways to get each total

OpenStudy (anonymous):

for example, there are 6 different ways to roll a total of 7 as a list they look like this: (1,6), (2,5), (3,4) , (4,3), (5,2), (6,1)

OpenStudy (anonymous):

there is only one way to roll a total of 2, namely (1,1) and also there is only one way to get a total of 12, namely (6,6)

OpenStudy (anonymous):

you should see the nice pattern as suggested by @dumbcow above

jimthompson5910 (jim_thompson5910):

alternative way: as dumbcow pointed out "well '7' is most likely, then its a symmetric distribution... 6 same as 8 5 same as 9 and so on..." there is a trick/shortcut: if the two numbers add to 14, then the two numbers are equally likely. Notice how 7+7 = 14, 6+8 = 14, 5+9 = 14, etc since 2+12 = 14, this means that 2 and 12 have the same probability of occurring

OpenStudy (emilyjones284):

Ohhhhh okay, I just figure it out right before I read @jim_thompson5910 's answer lol thanks for all your help both of you @satellite73